OpenStudy (anonymous):

Medal if solved!!!!!!!!!! Solve the system of equations below 5x - 2y+ 18 3x + 3y+ 15

OpenStudy (john_es):

May be is something like this. \[\begin{cases}5x-2y=18\\ 3x+3y=15\end{cases}\]?

OpenStudy (anonymous):

Yeah like that

OpenStudy (john_es):

There are a lot of ways to solve it. Buy first, I would simplify the second equation, dividing by 3, \[5x−2y=18\\x+y=5\]

OpenStudy (john_es):

Now you can use reduction, for example, or substitution.